If the measure of an angle is 30 less than twice the measure of its supplement, what is the measure of each angle?
Answer #1
x+3x-30=180 4x=180+30 4x=210 x=210/4 x=52.5 ans. for the smaller angle. 3*52.5-30=157.50-30=127.7 ans for the larger angle. Proof: 52.5+127.5=180 180=180
answered by: Stephanie24